There are 200 heads on a farm and 540 feet.
How many cows (4 legs) and how many chickens (2 legs) are there?

Answer by Edwin McCravy(20086)

There are 200 heads on a farm and 540 feet.
How many cows (4 legs) and how many chickens (2 legs) are there?
Two ways to do it. In your head and by algebra.

In your head:
540 feet means 270 pairs of feet, 200 of which are pairs of front feet,
since all animals have 1 head and 1 pair of front feet.  The remaining 70
pairs of feet are pairs of hind feet, which belong to the 70 cows, so the
other 130 animals are chickens.

By algebra:
x = number of cows
y = number of chickens

 x +  y = 200
4x + 2y = 540

Multiply the top equation by -2

-2x - 2y = -400
 4x + 2y = 540
--------------
 2x      = 140
       x = 70  

 x + y = 200
70 + y = 200
     y = 130.

Edwin